a gallon contains 4 quarts. A cartooning machine can fill 120 one-quart cartons a minute. How long will it take to put 600 gallons of orange juice into cartons?
Found 2 solutions by stanbon, ptaylor:
Answer by stanbon(75887)

a gallon contains 4 quarts. A cartooning machine can fill 120 one-quart cartons a minute. How long will it take to put 600 gallons of orange juice into cartons?
----------
120 one-quart cartons = 120/4 gallons = 30 gallons
If the machine produces 30 gallons per minute set up a proportion to
get the time required to produce 600 gallons:
x min/600 gal = 1min/30 gal
x=(600/30)(min)
x=20 minutes
===============
It will take 20 min for the machine to produce 600 gallons.
Cheers,
Stan H.

Answer by ptaylor(2198)

1 gal =4 qts
600 gal =4(600)=2400 qts
Machine fills 120 one-qt cartons in 1 minute
Then machine can fill 1200 one-qt cartons in 10 minutes
And the machine can fill 2400 one-qt cartons in 20 minutes---------ans
Another way to look at this problem:
Let x=time it takes to put 600 gal of orange juice into one-qt cartons
Now we know that x min is to 2400 qts as 1 min is to 120 qts:
We write this ratio as:
x/2400=1/120 multiply both sides by 2400
(x(2400))/2400 =2400/120 simplify
x=20 min
Hope this helps----ptaylor